# A Multi-Messenger Search for the Supermassive Black Hole Binary in 3C 66B with the Parkes Pulsar Timing Array

### v1

@@ 3.2 Bayesian inference
where vector $\bm{\mu}$, a function of $\bm{\theta}$, represents a prediction of deterministic contributions to pulse arrival times $\bm{\delta t}$, which includes a CGW model, as well as deterministic noise transients such as exponential dips in pulse arrival times (Goncharov et al., 2021b). The form of our covariance matrix $\bm{C}$ and the methods for computing it are described in detail in Arzoumanian et al. (2016). In short, $\bm{C}=\bm{N}+\bm{T}^{\text{T}}\bm{B}^{-1}\bm{T}$, where diagonal matrix $\bm{N}$ represents temporally uncorrelated “white” noise, $\bm{T}$ is the design matrix for reduced-rank modeling (Lentati et al., 2013), and $\bm{B}$ is the covariance matrix of reduced-rank coefficients $\bm{b}$. The likelihood is marginalised over $\bm{b}$, which determines the time series realisation $\bm{T}\bm{b}$. Thus, the design matrix $\bm{T}$ is a mapping from the parameter space $\bm{b}$ to the time series. It is made up of three blocks. Block $\bm{M}$ maps timing model parameters to the time series. Block $\bm{F}$ maps Fourier sine and cosine amplitudes of time-correlated “red” noise to the time series. Block $\bm{U}$ maps delays in arrival times at different observing epochs due to pulse jitter to all arrival times. Coefficients $\bm{b}$ are assumed to be samples from the zero-mean Gaussian distribution described by the covariance matrix $\bm{B}$. For the coefficients of the pulsar timing model, we assume an uninformative prior with a standard deviation of $10^{40}$ in respective units, which is much greater than the typical coefficient value.
3.3 Noise Model
@@
Noise modeling in PTAs is crucial for correct astrophysical interpretability of the gravitational wave searches (Goncharov et al., 2024). Broadly, modeled PTA noise is categorized into temporally-uncorrelated (white) noise, temporally-correlated (red) noise with the power spectral density modeled as a power law, as well as noise transients. We use the noise model from Reardon et al. (2023). In addition to this, we model the contribution of the common-spectrum noise process associated with the GWB (Goncharov et al., 2022). It is shown in Zhao et al. (2025) that not modeling this process leads to false positives. Agazie et al. (2023b) suggests that further modeling of the Hellings & Downs (1983) correlations may be necessary for the data where they are evident. This may be required for future CGW searches with the PPTA data. Furthermore, we impose a minimum allowed value of the red noise power spectral density of $T_{\text{obs}}\times 10^{-20}~[\text{s}^{3}]$, which is well below the noise in PPTA DR3. This is done to avoid numerical errors in the likelihood.
3.4 Analysis of the 3-mm data
@@
To perform an re-analysis of the 3C 66B 3-mm flux measurements from Iguchi et al. (2010), we employ the univariate Gaussian likelihood, $\mathcal{L}(\bm{S}|\bm{\theta}_{\text{EM}})$, where $\bm{S}$ is a vector of flux measurements in units mJy. The general form is the same as Equation 4, except that the covariance matrix is diagonal with elements $\bm{\sigma}_{\text{EM}}^{2}=\bm{\sigma}_{S}^{2}e_{\text{f}}^{2}+e_{\text{q}}^{2}$. Here, $\bm{\sigma}_{S}$ is the flux measurement uncertainty. White noise parameters $(e_{\text{f}},e_{\text{q}})$ are the error factor and the error added in quadrature. The analogue of $\bm{\mu}$ from Equation 4 for the EM data is the sine-cubic function from Iguchi et al. (2010):
@@
The chirp mass is calculated from $\bm{\theta}_{\text{EM}}$ and Equation 5 as follows. Iguchi et al. (2010) suggest that the sine component is due to the Doppler boosting of the flux from the core of 3C 66B due to binary motion around the center of mass, whereas the cubic component may be responsible for the non-thermal radiation in the knots of the jet in the core. From the frequency of the sine component $f_{\text{EM}}$, the SMBHB orbital frequency $f_{\text{k}}$ from Sudou et al. (2003), and the apparent boost in the relativistic jet $\beta_{\text{app}}$ (Iguchi et al., 2010), we compute the jet outflow boost $\beta$ and the inclination angle $\iota$ between the jet axis and the line of sight. The flux ratio from Doppler boosting is $\bm{S}_{\text{r}}=\max(\bm{S})/\min(\bm{S})$, where $\max(\bm{S})=\bar{S}+A_{\text{EM}}$ and $\min(\bm{S})=\bar{S}-A_{\text{EM}}$ are computed from the mean flux $\bar{S}$. From the above, assuming the power law relation between the Doppler factor and $\bm{S}_{\text{r}}$ with a spectral index $\alpha$, we calculate the SMBHB orbital radius. Considering that the jet originates from one of the black holes in a binary, but not two, and from the total mass of the binary $M_{\text{tot}}$, which can be computed based on the galaxy velocity dispersion relation, a chirp mass $\mathcal{M}$ is obtained. To sum up, $\mathcal{M}=\mathcal{G}(\theta_{\text{EM}},\bm{\varphi})$, where $\bm{\varphi}=(f_{\text{k}}$, $\beta_{\text{app}}$, $M_{\text{tot}}$, $\alpha)$ and $\mathcal{G}$ is a function with the aforementioned operations. For more details, please refer to Appendix B and Iguchi et al. (2010).
3.5 Multi-messenger likelihood
Here, we lay out the groundwork for the joint analysis of CGW and EM data, which may be explored in future work. Because EM and CGW data are independent, the total likelihood of these data is a product of the likelihoods: $\mathcal{L}(\bm{\delta t},\bm{S}|\bm{\theta},\bm{\theta}_{\text{EM}})=\mathcal{L}(\bm{\delta t}|\bm{\theta})\mathcal{L}(\bm{S}|\bm{\theta}_{\text{EM}})$, where $\bm{\theta}=(\bm{\theta}_{\text{GW}},\mathcal{M})$. Effectively, the common parameter in both the CGW and the EM model from Iguchi et al. (2010) is the chirp mass $\mathcal{M}$. Therefore, the posterior is $\mathcal{P}(\bm{\theta}_{\text{GW}},\bm{\theta}_{\text{EM}},\varphi|\bm{\delta t},\bm{S})=\mathcal{L}(\bm{\delta t},\bm{S}|\bm{\theta}_{\text{GW}},\bm{\theta}_{\text{EM}},\varphi)\pi(\bm{\theta}_{\text{GW}})\pi(\bm{\theta}_{\text{EM}})\pi(\bm{\varphi})$. Here, we also express $\mathcal{L}(\bm{\delta t}|\bm{\theta}_{\text{GW}},\theta_{\text{EM}},\varphi)=\mathcal{L}(\bm{\delta t}|\bm{\theta}_{\text{GW}},\mathcal{M}=\mathcal{G}(\theta_{\text{EM}},\bm{\varphi}))$.
4 Results
4.1 Model selection
@@
4.2 Limits on the SMBHB chirp mass at 3C 66B
@@
We perform parameter estimation of the log-10 chirp mass. As in the previous calculation, we assume the luminosity distance as a fixed parameter based on the known redshift of 3C 66B and Planck 2018 results (Planck Collaboration et al., 2020). The difference is that now we use a broad uniform prior $\pi(\log_{10}\mathcal{M})=\mathcal{U}(7.5,9.5)~[\log_{10}M_{\odot}]$. This provides a more EM-independent perspective consistent with the decreasing probability of finding an SMBHB towards high $\mathcal{M}>10^{9}~M_{\odot}$ in nature at $z=0$ (Hopkins et al., 2007), defined by the SMBHB mass function. We also use a more convenient prior on $\pi(\log_{10}(f_{\text{GW}}))=\mathcal{N}(-7.22,0.01)~[\log_{10}\text{Hz}]$, which is still consistent with the EM model.

### v2

@@ III.2 Bayesian Inference
where vector $\bm{\mu}$, a function of $\bm{\theta}$, represents a prediction of deterministic contributions to pulse arrival times $\bm{\delta t}$, which includes a CGW model, as well as deterministic noise transients such as exponential dips in the pulse arrival times (Goncharov et al., 2021a). The form of our covariance matrix $\bm{C}$ and the methods for computing it are described in detail in Arzoumanian et al. (2016). In short, $\bm{C}=\bm{N}+\bm{T}\bm{B}\bm{T}^{\text{T}}$, where the block diagonal matrix $\bm{N}$ represents temporally uncorrelated “white” noise, $\bm{T}$ is the design matrix for reduced-rank modeling (Lentati et al., 2013), and $\bm{B}$ is the prior covariance matrix of reduced-rank coefficients $\bm{b}$. The likelihood is marginalized over $\bm{b}$, which determines the time series realization $\bm{T}\bm{b}$. Thus, the design matrix $\bm{T}$ is a mapping from the parameter space $\bm{b}$ to the time series. It is made up of three blocks. Block $\bm{M}$ maps timing model parameters to the time series. Block $\bm{F}$ maps Fourier sine and cosine amplitudes of time-correlated “red” noise to the time series. Block $\bm{U}$ maps delays in the arrival times at different observing epochs due to pulse jitter to all arrival times. Coefficients $\bm{b}$ are assumed to be samples from the zero-mean Gaussian distribution described by the covariance matrix $\bm{B}$. For the coefficients of the pulsar timing model, we assume an uninformative prior with a standard deviation of $10^{40}$ in respective units, which is much greater than the typical coefficient value.
III.3 Noise Model
@@
Noise modeling in PTAs is crucial for correct astrophysical interpretability of the GW searches (Goncharov and Sardana, 2025; Hazboun et al., 2020; Larsen et al., 2024; Reardon et al., 2023). Broadly, modeled PTA noise is categorized into temporally uncorrelated (white) noise, temporally correlated (red) noise with the power spectral density modeled as a power law, as well as noise transients. We use the noise model from Reardon et al. (2023). In addition to this, we model the contribution of the common-spectrum noise process associated with the GWB (Goncharov et al., 2022). It is shown in Zhao et al. (2025) that not modeling this process leads to false positives. Agazie et al. (2023b) suggest that further modeling of the Hellings and Downs (1983) correlations may be necessary for the data where they are evident. We do not model these correlations in this work; however, it may be required for future CGW searches with the PPTA data. Furthermore, we impose a minimum allowed value of the red noise power spectral density of $T_{\text{obs}}\times 10^{-20}~\text{s}^{3}$, which is well below the noise in PPTA DR3. This is done to avoid numerical errors in the likelihood.
III.4 Analysis of the 3 mm Data
@@
We perform a Bayesian reanalysis of the 3C 66B 3 mm flux measurements rather than using the original frequentist results from Iguchi et al. (2010). This allows a direct comparison between the GW posteriors and the EM posterior. We employ the univariate Gaussian likelihood, $\mathcal{L}(\bm{S}|\bm{\theta}_{\text{EM}})$, where $\bm{S}$ is a vector of flux measurements in units of mJy. The general form is the same as Equation 4, except that the covariance matrix is diagonal with elements $\bm{\sigma}_{\text{EM}}^{2}=\bm{\sigma}_{S}^{2}e_{\text{f}}^{2}+e_{\text{q}}^{2}$. Here, $\bm{\sigma}_{S}$ is the flux measurement uncertainty. White noise parameters $(e_{\text{f}},e_{\text{q}})$ are the error factor and the error added in quadrature, respectively. The analog of $\bm{\mu}$ from Equation 4 for the EM data is the sine cubic function from Iguchi et al. (2010):
@@
The chirp mass is calculated from $\bm{\theta}_{\text{EM}}$ and Equation 5 as follows. Iguchi et al. (2010) suggest that the sine component is due to the Doppler boosting of the flux from the core of 3C 66B due to binary motion around the center of mass, whereas the cubic component may be responsible for the nonthermal radiation in the knots of the jet in the core. From the frequency of the sine component $f_{\text{EM}}$, the SMBHB orbital frequency $f_{\text{k}}$ from Sudou et al. (2003), and the apparent boost in the relativistic jet $\beta_{\text{app}}$ (Iguchi et al., 2010), we compute the jet outflow boost $\beta$ and the inclination angle $\iota$ between the jet axis and the line of sight. The flux ratio from Doppler boosting (Rieger and Mannheim, 2000; De Paolis et al., 2002, 2004) is $\bm{S}_{\text{r}}=\max(\bm{S})/\min(\bm{S})$, where $\max(\bm{S})=\bar{S}+A_{\text{EM}}$ and $\min(\bm{S})=\bar{S}-A_{\text{EM}}$ are computed from the mean flux $\bar{S}$. From the above, assuming a power-law relation between the Doppler factor and $\bm{S}_{\text{r}}$ with spectral index $\alpha$, we calculate the SMBHB orbital radius. Considering that the jet originates from one of the black holes in a binary, but not two, and from the total mass of the binary $M_{\text{tot}}$, which can be computed based on the galaxy velocity dispersion relation, the chirp mass $\mathcal{M}$ is obtained. To sum up, $\mathcal{M}=\mathcal{G}(\theta_{\text{EM}},\bm{\varphi})$, where $\bm{\varphi}=(f_{\text{k}}$, $\beta_{\text{app}}$, $M_{\text{tot}}$, $\alpha)$ and $\mathcal{G}$ is a function with the aforementioned operations. For more details, please refer to Appendix B and Iguchi et al. (2010).
IV Results
IV.1 Model Selection
@@
Exponentiating our result, the Bayes factor is $\mathcal{B}=0.9973(6)$, whereas Agarwal et al. (2025) find a value of $\mathcal{B}=0.293(3)$. This disagreement is due to noise fluctuations, which are capable of yielding Bayes factors of up to $\sim 1$ according to the null distribution (see Figure 1, left, in Agarwal et al., 2025). The use of different pulsars, different instruments, as well as different noise models can influence the fluctuations. Additionally, the Bayes factor calculated in this study is based on the Earth term, whereas the Bayes factor in Agarwal et al. (2025) is calculated based on both the Earth term and the pulsar term.
IV.2 Limits on the Supermassive Black Hole Binary Chirp Mass in 3C 66B
@@
We perform parameter estimation of the log-10 chirp mass. As in the previous calculation, we assume the luminosity distance as a fixed parameter based on the known redshift of 3C 66B and Planck 2018 results (Planck Collaboration et al., 2020). The difference is that now we use a broad uniform prior $\pi(\log_{10}\mathcal{M})=\mathcal{U}(7.5,9.5)~[\log_{10}M_{\odot}]$. This provides a more EM-independent perspective consistent with the decreasing probability of finding an SMBHB toward high $\mathcal{M}>10^{9}~M_{\odot}$ in nature at $z=0$ (Hopkins et al., 2007), defined by the SMBHB mass function. We also use a more convenient prior on $\pi(\log_{10}(f_{\text{GW}}))=\mathcal{N}(-7.22,0.01)~[\log_{10}\text{Hz}]$, which is still consistent with the EM model.
The results are shown in Figure 2. When only the Earth term is modeled in our analysis, we find an upper limit of $\mathcal{M}<8.25\times 10^{8}~M_{\odot}$ at 95% credibility. When including the pulsar term, we find a more constraining limit, $\mathcal{M}<6.90\times 10^{8}~M_{\odot}$. For comparison, we show the value from the EM model from Iguchi et al. (2010), $(7.9^{+3.8}_{-4.5})\times 10^{8}$ $M_{\odot}$. Our limits are not below the prediction of the EM model. Therefore, we cannot rule it out. However, the tightest upper limit calculated rules out 58% of the uncertainty region of the EM-model. For consistency with Arzoumanian et al. (2023) and Agarwal et al. (2025), we also obtain a posterior by reweighting our posteriors to a uniform prior in linear space $\pi(\mathcal{M})=\mathcal{U}(10^{7.5},10^{9.5})~[M_{\odot}]$. While not matching astrophysical expectations for the SMBHB mass function, this presents the most conservative limits. With this prior, we report an upper limit at $95\%$ credibility to be $\mathcal{M}<11.06\times 10^{8}~M_{\odot}$ for the Earth term and $\mathcal{M}<10.63\ \times 10^{8}~M_{\odot}$ when the pulsar term is included. We find a 16% improvement in the chirp mass upper limit when including the pulsar term. This difference is on the larger end of the 5-15% improvement found in previous studies (Charisi et al., 2024; Zhu et al., 2016), which compare cases where the signal is detected with a signal-to-noise ratio between 5 and 15. Upper limits use information from the tail of the distribution, whereas detections use the bulk of the distribution, which may explain the larger improvement when including the pulsar term in this work. Our limits are a factor of 1.55 lower compared to Arzoumanian et al. (2020b), a factor of 1.33 lower compared to Arzoumanian et al. (2023), and similar to Agarwal et al. (2025).
We also report the result of our reanalysis of the 3 mm data from Iguchi et al. (2010). Following the methodology in Section III.4, we obtain a posterior on the log-10 chirp mass of the SMBHB in 3C 66B. It is shown as the black shaded area in Figure 2. The posterior is skewed toward higher masses, consistently with the uncertainties reported in Iguchi et al. (2010) and presented in Table 1.
Figure 2: Observational constraints on the chirp mass $\mathcal{M}$ of an SMBHB in 3C 66B. The histograms show a posterior for $\mathcal{M}$ based on the analysis of PPTA DR3. The dashed lines show the respective upper limits at $95\%$ credibility based on the astrophysically motivated prior $\pi(\log_{10}\mathcal{M})=\mathcal{U}(7.5,9.5)~[\log_{10}M_{\odot}]$. The dotted lines show the limits based on the conservative prior $\pi(\mathcal{M})=\mathcal{U}(10^{7.5},10^{9.5})~[M_{\odot}]$. In green, we show the results of the analysis with only the Earth term of the signal modeled. In blue, we show the results based on modeling both the Earth term and the pulsar term of the CGW. The black solid line with arbitrary density normalization represents the posterior we obtain from the 3 mm data analyzed in Iguchi et al. (2010). The vertical orange line and the band correspond to the value and the uncertainty reported in Iguchi et al. (2010).
IV.3 Limits on the Gravitational-wave Strain from 3C 66B
Instead of calculating the strain amplitude $h_{0}$ of the CGW from 3C 66B based on the luminosity distance using Equation 2, we now treat it as a free parameter. This way, the distance to 3C 66B is not involved in the calculation. Similar to the chirp mass prior, we use the prior $\pi(\log_{10}h_{0})=\mathcal{U}(-17.0,-12.5)$, which is more consistent with the decreasing probability of an SMBHB with a high $h_{0}$. The other priors are as in Section IV.2. The posterior on $h_{0}$ is shown in Figure 3. With only the Earth term, we find an upper limit at 95% credibility to be $\log_{10}h_{0}<-14.12$. When including the pulsar term, we find a more constraining limit, $\log_{10}h_{0}<-14.44$. Using the more conservative prior, $\pi(h_{0})=\mathcal{U}(10^{-17.0},10^{-12.5})$, we find an upper limit for the Earth term of $\log_{10}h_{0}<-13.86$ and an upper limit of $\log_{10}h_{0}<-14.20$ when including the pulsar term. As in Section IV.2, the range of $h_{0}$ of the EM model from Iguchi et al. (2010) is partially ruled out in all cases.
It is also of interest to compare the strain amplitude of a CGW signal from 3C 66B with the expectation for the GWB at the EM-predicted frequency of the binary. Goncharov et al. (2025) propose a fiducial prior on the characteristic strain amplitude of the background, $h_{\text{c}}(f)=A(f~\text{yr})^{-2/3}$, characteristic of SMBHBs. Where $A$ is the strain amplitude at the gravitational wave frequency $f$. The prior is based on black hole masses inferred from kinematic observations and galaxy velocity dispersions. Based on this prior, we calculate $h_{0,\text{GWB}}=2h_{\text{c}}/(2^{2/3}\sqrt{32/5})$, to revert the rms averaging over the SMBHB inclination and polarization assumed for the background (see Equation 8 in Burke-Spolaor et al. (2019)). The best-fit $h_{\text{c}}(f=\text{yr}^{-1})$ of the GWB from the data of the European Pulsar Timing Array by Goncharov et al. (2025) is within $3\sigma$ interval of the prior, above the prior mean. Thus, the prior is consistent with observations.
  • the EM model of 3C 66B corresponds to a loud source in the expected GWB;
  • a fraction of the parameter space of the EM model for 3C 66B, which we have ruled out, is also not supported by the predicted GWB amplitude; and
  • our limits on $f_{\text{CGW}}$ of 3C 66B are, on average, higher than the expected GWB because we are limited by white noise at the frequency of 3C 66B.
The value of $h_{0,\text{GWB}}$ can also be interpreted as the equivalent expected CGW amplitude of the background if it is dominated by one source (3C 66B) at the 3C 66B CGW frequency. This scenario is possible, given the scaling of the number of SMBHB sources contributing to the background with frequency. This may also lead to a deviation from the power-law model for $h_{\text{c}}(f)$ at the frequency of 3C 66B, depending on the number density of SMBHBs (Phinney, 2001; Agazie et al., 2025). It is also worth noting that Zhu et al. (2019) find that the EM model for 3C 66B is in tension with the observed GWB amplitude based on the expected SMBHB merger rate.
Figure 3: Observational constraints on the CGW strain amplitude, $h_{0}$, of an SMBHB in 3C 66B. The histograms show the posterior on $h_{0}$ based on the analysis of PPTA DR3. The dashed lines show the respective upper limits at $95\%$ credibility based on the astrophysically motivated prior $\pi(\log_{10}h_{0})=\mathcal{U}(-17.0,-12.5)$. The dotted lines show the limits based on the conservative prior $\pi(h_{0})=\mathcal{U}(10^{-17.0},10^{-12.5})$. In green, we show the results of the analysis with only the Earth term of the signal modeled. In blue, we show the results based on modeling both the Earth term and the pulsar term of the CGW. The black solid line with arbitrary density normalization represents the theoretically expected GWB contribution at the frequency of 3C 66B. The vertical orange line and the band correspond to the value and the uncertainty reported in Iguchi et al. (2010).
IV.4 Upper Limits as a Function of the Continuous Gravitational-Wave Frequency
Following Arzoumanian et al. (2020b), for a broader perspective on how PPTA DR3 limits on $\mathcal{M}$ depend on $f_{\text{GW}}$ for any SMBHB at the position of 3C 66B, we show Figure 4. For this calculation, we use the same prior as in Section IV.2, except the prior on the CGW frequency is now $\pi(\log_{10}f_{\text{GW}})=\mathcal{U}(-9.0,-7.0)~[\log_{10}\text{Hz}]$. We also only model the Earth term of the CGW signal. The figure highlights that the PPTA is well suited to target 3C 66B, as the value suggested by the EM model is at the frequency to which our targeted search is most sensitive. Consistent with Figure 2, when modeling the pulsar term and using astrophysically motivated priors, we are close to ruling out the best-fit value of the EM model. We observe the highest density of posterior samples at around $f_{\text{GW}}=\text{yr}^{-1}$, which corresponds to uncertainties associated with measuring pulsar positions. There is also a minor overdensity of posterior samples at $f_{\text{GW}}=(2~\text{yr})^{-1}$, which corresponds to uncertainties associated with measuring pulsar parallaxes. Given that the posterior is not uniform at these frequencies, some of the uncertainties may be systematic. We find that the chirp mass upper limits are constrained to lower values in the targeted search as compared to the results from Figure 7 in the PPTA DR3 all-sky search (Zhao et al., 2025).
Figure 4: Posterior for the chirp mass $\mathcal{M}$ and frequency $f_{\text{GW}}$ of 3C 66B. The color map shows the number of posterior samples in $\mathcal{M}$-$f_{\text{GW}}$ space. The best-fit value from Iguchi et al. (2010) is shown as a red star. The blue solid line shows the respective upper limit at $95\%$ credibility based on the astrophysically motivated prior $\pi(\log_{10}\mathcal{M})=\mathcal{U}(7.5,10.5)~[\log_{10}M_{\odot}]$. The yellow dashed line shows the limits based on the conservative prior $\pi(\mathcal{M})=\mathcal{U}(10^{7.5},10^{10.5})~[M_{\odot}]$.
IV.5 Frequentist Search
To confirm the nondetection of a CGW, we perform a frequentist targeted search based on the $\mathcal{F}$ statistic (Ellis et al., 2012), a frequentist measure of evidence for a CGW in our data. In particular, we chose to calculate the $\mathcal{F}_{\text{e}}$ statistic, which is based on only modeling the Earth term of the CGW signal. This statistic is a function of CGW frequency and sky position, which we set to the 3C 66B EM model values. We find a value of $\mathcal{F}_{\text{e}}=4.90$, corresponding to the false alarm probability $\text{FAP}=0.50$. Ellis et al. (2012) suggest a threshold for FAP to be $10^{-4}$ to claim a detection. The trials factor for our search is one. The background distribution of the $\mathcal{F}_{\text{e}}$ statistic is the chi-squared distribution with four degrees of freedom and is displayed in Figure 5 along with the value of the $2\mathcal{F}_{\text{e}}$ statistic. Based on this null distribution, we find a $p\text{-value}=0.04$ which corresponds to about $2.1\sigma$. Thus, the $\mathcal{F}_{\text{e}}$ statistic value we find is consistent with noise.
Figure 5: The background distribution of the $\mathcal{F}_{\text{e}}$ statistic, which is a chi-squared distribution with four degrees of freedom. The green vertical dashed line is the $2\mathcal{F}_{\text{e}}$ statistic calculated at the frequency and sky position of 3C 66B.
V Discussion
V.1 Multimessenger Joint Likelihood
Here, we lay out the groundwork for the joint analysis of CGW and EM data, which may be explored in future work. Because EM and CGW data are independent, the total likelihood of these data is a product of the likelihoods: $\mathcal{L}(\bm{\delta t},\bm{S}|\bm{\theta},\bm{\theta}_{\text{EM}})=\mathcal{L}(\bm{\delta t}|\bm{\theta})\mathcal{L}(\bm{S}|\bm{\theta}_{\text{EM}})$, where $\bm{\theta}=(\bm{\theta}_{\text{GW}},\mathcal{M})$. Effectively, the common parameter in both the CGW model and the EM model from Iguchi et al. (2010) is the chirp mass $\mathcal{M}$. Therefore, the posterior is $\mathcal{P}(\bm{\theta}_{\text{GW}},\bm{\theta}_{\text{EM}},\varphi|\bm{\delta t},\bm{S})=\mathcal{L}(\bm{\delta t},\bm{S}|\bm{\theta}_{\text{GW}},\bm{\theta}_{\text{EM}},\varphi)\pi(\bm{\theta}_{\text{GW}})\pi(\bm{\theta}_{\text{EM}})\pi(\bm{\varphi})$. Here, we also express $\mathcal{L}(\bm{\delta t}|\bm{\theta}_{\text{GW}},\theta_{\text{EM}},\varphi)=\mathcal{L}(\bm{\delta t}|\bm{\theta}_{\text{GW}},\mathcal{M})=\mathcal{G}(\theta_{\text{EM}},\bm{\varphi})$.
V.2 Feasibility of Standard Siren Cosmology using Targeted Searches with Pulsar Timing Arrays
Standard sirens are compact binaries that emit GW and EM radiation simultaneously, allowing for measurement of the Universe’s expansion rate (Schutz, 1986). The redshift of the host galaxy of a standard siren from EM information enables the determination of its recessional velocity $v$, whereas the proper distance $D$ is measured based on the GW signal. The Hubble–Lemaître constant, which quantifies the Universe’s expansion at the present time, is
$H_{0}=\frac{\nu}{D}.$ (6)